# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 4 Analysis
Let $z=|e^{*}|$. Then we know $\mathbb{E}[{z}]=\mu$ and by Markov inequality, $Pr[z>\mu/\epsilon]<\epsilon$, meaning that with probability $1-\epsilon$, the size of the infected set is at most $\mu/\epsilon$. Now we can modify Algorithm 1 to exploit this fact. The revised algorithm maintains the same execution as Algorithm 1, however, before initiation, it removes all the edges that have a size greater than $\mu/\epsilon$.
We know that with probability $1-\epsilon$, the realized edge is not removed and the largest edge has size at most $\mu/\epsilon$. Therefore, with probability $1-\epsilon$, the algorithm returns the correct $e^{*}$. The number of tests is argued as before, with the difference that the size of $\tilde{\mu}\leq\mu/\epsilon$ (because the largest edge has size at most $\mu/\epsilon$), so in the second round it performs at most $L_{2}\leq\frac{\mu}{\epsilon(1-2c)}$ tests and the proof is completed. ∎
@@
Markov bound works for general distributions, but it might happen that the mass in $\mathcal{D}$ is concentrated around the mean. In this case, the dependency on $1/\epsilon$ might be slower. For example, if the concentration is exponential, i.e. $Pr(|e^{*}|>2\mu)<2^{-\Omega(n)}$, then with probability $1-2^{-\Omega(n)}$ the algorithm recovers $e^{*}$ and the expected number of tests is $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{2\mu}{(1-2c)}$. More precisely, we have the following theorem.
@@
Suppose $e^{*}$ is sampled from the probability distribution $\mathcal{D}$ and we have $Pr(|e^{*}|>c\mu)=o(1)$, for a constant $c$. Then, with probability $1-o(1)$, Algorithm 1 returns the correct edge $e^{*}$ using $O(H(X)+\mu)$ tests with probability $1-o(1)$.
@@ 5 Recovering/Improving Prior Results
In [NSG+21a], as we discussed briefly in Section 2.2, the authors consider $F$ families where each family $j$ has size $M_{j}$, so $n=\sum_{j}M_{j}$. Each family $j$ is infected with probability $q$ and when a family is infected, all of its nodes are infected with probability $p_{j}$ independent of the rest. If a family is not infected, then all its nodes are not infected. They show that to have a constant probability of recovery, one needs $L\geq H(X)=Fh_{2}(q)+\sum_{j=1}^{F}qM_{j}h_{2}\left(p_{j}\right)-w_{j}h_{2}\left(\frac{1-q}{w_{j}}\right)$ where $w_{j}=1-q+q\left(1-p_{j}\right)^{M_{j}}$ and $h_{2}$ is the binary entropy. For the symmetric case where $\forall jp_{j}=p,M_{j}=M$ and hence $w_{j}=w$, the entropy simplifies to $H(X)=Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})$. They provide an algorithm that in expectation performs at most $O(Fq(\log(F)+M)+nqp\cdot\log n)$.
Note that $\mu=nqp$, and by Theorem 4.10, with probability $1-\epsilon$ the target edge can be found using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+\frac{nqp}{\epsilon})$ tests for a constant $c$. If $F\gg 1$ and $M\gg 1$, then the expected number of infections is concentrated around its mean, and with probability $1-o(1)$ the algorithm succeeds using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+{nqp})$. This significantly improves the upper bound provided in [NSG+21a].
@@
In [ACÖ23], the authors consider a stochastic block infection model (SBIM) where there are $m$ communities and each one has $k$ nodes, so $n=mk$. Each node is a seed with probability $p$ independently. Then every seed node infects nodes in the same community with probability $q_{1}$ and $q_{2}$ for the nodes outside of the community, $q_{1}>q_{2}$. The authors show lower and upper bounds on the number of tests. For the special case where $q_{2}=0$, they further show optimality of the proposed algorithm. Consider $q_{2}=0$ and let $q=q_{1}$. In this case, when $kp\ll 1\text{, and }q\ll\frac{1}{\sqrt{k\log\left(\frac{1}{kp}\right)}},$ the entropy is given by $H(X)=m\cdot k\cdot p\cdot\log\left(\frac{1}{kp}\right)+m\cdot k^{2}\cdot p\cdot q\cdot\left(\log k+\log\log\left(\frac{1}{kp}\right)\right)+1$. The authors show that their algorithm is order-optimal when, in addition to the above two conditions $kp\ll 1$ and $q\ll\frac{1}{\sqrt{k\log(\frac{1}{kp})}}$, we have $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$.
In Appendix D, we show that with only assuming $kp\ll m^{-\beta}$, the conditions of Corollary 4.11 hold and, the infected set is recovered with $O(H(X)+\mu)$ tests in expectation. In [ACÖ23], the authors prove $\mu\leq m\cdot k^{2}\cdot p\cdot(1/k+q)$. It is not hard to show that when $1/k\leq q$, which is a relaxed condition compared to $1\ll kq$, we have $\mu\leq H(X)$ and hence the dominating term in $O(H(X)+\mu)$ is $H(X)$, i.e. $O(H(X)+\mu)=O(H(X))$ and Corollary 4.11 thus provides an order-optimal performance guarantee. More generally, Theorem 4.10 provides the following result: Algorithm 1 can recover the infected set with probability $1-\epsilon$ using at most $O(H(X)+\mu/\epsilon)$ tests in expectation. This allows one to relax two additional conditions that [ACÖ23] has imposed, i.e. $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$.
@@ 6 Lower Bound and Optimality of Algorithm 1
In the probabilistic group testing literature, entropy $H(X)$ constitutes an order-wise tight lower bound [LCH+14] (see also Theorem 2.2.) The upper bounds we established in Theorems 4.8-4.10 and Corollary 4.11 involve not only $H(X)$ but also $\mu$, which is the expected number of infections. To shed light on the extent to which our proposed algorithm is order-optimal and our analytical upper bounds are tight, we seek to answer the following questions. (i) Is entropy always a tight lower bound in the generalized group testing problem with general correlation? (ii) Does $\mu$ play a fundamental role (lower-bounding the number of tests) or should one hope to tighten it in our results? (iii) Is the upper bound established in Theorem 4.10 and/or its corresponding algorithm order-wise optimal?
@@ 6.1 When Entropy is a Loose Lower Bound
Graph $G=(V,E)$ has $n$ nodes where each edge of size $n-1$ is included in $E$. To each each edge $e$, assign a uniform probability mass $p(e)=\frac{1}{n}$. Figure 2 demonstrates such a graph with $n=4$.
In the above example, the entropy is $\log n$. What makes this example important for us is that any algorithm with a constant probability of success requires at least $\Omega(n)$ tests. This can be proved using a very similar argument as in [GLS22, Theorem 4], and for the sake of completeness, we provide a short proof in Appendix B. What is significant here is that the gap between the entropy and the required number of tests is $O(n/\log n)$, which is much larger than expected. We can generalize the above example as follows:
@@
Graph $G=(V,E)$ has $n$ nodes where at least $c$ fraction of all edges of size $d$ is included in $E$, for a constant $c\in(0,1)$. a uniform probability mass $p(e)=\frac{1}{|E|}$ is assigned to each edge $e$.
In Appendix C, we prove that any algorithm performing on Example 6.2 and achieving a constant probability of success needs to perform $\Omega(d)$ tests. These examples raise the question of finding other lower bounds or identifying classes of graphs where entropy is not a tight lower bound. Before proceeding further, suppose we run Algorithm 1 on this example. It is worth noting that for any constant $c$, the algorithm bypasses the initial stage and proceeds to test nodes individually in line 1, requiring $n$ tests. This example highlights scenarios where conducting individual tests is optimal, leaving no room for improving the theoretical guarantees in Theorem 4.10 and the following corollaries.
@@
Now, we show that the Example 6.1 is not the only graph where the entropy is loose, and we generalize it for a larger family of graphs where $H(X)$ is a loose lower bound and at least $\mu$ tests are needed. Consider $n=|V|$, and let $m$, $d=|e|$ be given parameters. We build a graph with $n$ nodes and $|E|=m+O(d)$ edges where each edge has size $d$. Let the distribution $\mathcal{D}$ be uniform over the edges. Here, $H(X)=\log(m+O(d))$ and $\mu=d$. The graph is constructed in the following manner, where we start with an empty set for the edges and gradually add edges:
@@
The first time that $|E|$ exceeds $m$, there are at most $d$ edges added, hence $|E|=m+O(d)$. We claim that at least $\Omega(d)$ tests are needed to recover the infected set. Suppose $e^{*}$ is the sampled edge, so at some point in line 2 $e^{*}$ was added to $E$. Hence there is a set $S$ of size $d+1$ that contains $e^{*}$ and there are $d$ edges inside of $S$. Now even if the algorithm knows the set $S$, using a similar argument as in [GLS22] and Appendix B, the algorithm still needs $\Omega(d)$ tests as only one node is negative and it is equally likely to be any of them.
Now if $\mu=d>\log(m+O(d))=H(X)$, then the entropy lower bound is loose and $\mu$ is an order-wise tight lower bound. This happens when $2^{d}>m+O(d)$.
@@ 6.2 When $\mu$ is not a Lower Bound
Consider a graph $G=(V,E)$ comprising $n$ nodes and the edges are of the form $e_{i}=\{v_{1},\ldots,v_{i}\}$ and $p_{e_{i}}=1/n$ for each $i$.
@@
Consider a graph with vertex set $V=V_{1}\cup V_{2}\cdots\cup V_{n}$ where $|V_{i}|=n$, so $|V|=n^{2}$ and we call each $V_{i}$ a community. Community $V_{i}$ is composed of nodes $V_{i,j}$ for $j=1,\ldots,n$. The edge set is composed of “small edges” and “large edges”. The set of small edges is defined by $E_{1}=\{e\mid\forall j,k:e=V_{j}\setminus\{V_{j,k}\}\}$. So there are $n^{2}$ small edges and each small edge has size $n-1$. The set of large edges is defined by $E_{2}=\{e\mid\forall i:e=V\setminus V_{i}\}$, so there are $n$ of them and each has size $n(n-1)$. Here, $E=E_{1}\cup E_{2}$. The distribution $\mathcal{D}$ puts half of the mass on $E_{1}$ and half of the mass on $E_{2}$, and the edges in $E_{1}$ (respectively $E_{2}$) have the same probability among themselves. Figure 6 shows and example with $n=4$.
Figure 6: The graph in Example 6.5 with $n=4$, 16 nodes, and 20 edges. “Small” edges have black border, and “large” edges have red border.
In the above example, $\mu=\Omega(n^{2})$ and $H(X)=O(\log n)$. Following Example 6.1, it is not hard to see that $\Omega(n)$ tests are needed, because even after knowing whether the sampled edge is in $E_{1}$ or $E_{2}$, the algorithm needs to find the single negative node (in some community $V_{i}$) to recover the sampled edge.
But $O(n)$ tests are sufficient: one can pick a single node from each $V_{1},V_{2},V_{3}$ and test them individually. If at least two of them are negative, it means the sampled edge is in $E_{1}$, otherwise, it is in $E_{2}$. If the sampled edge is in $E_{1}$, using $\log n$ tests the infected $V_{i}$ can be detected and by testing all of its nodes, the sampled edge is recovered. If the sampled edge is in $E_{2}$, we can find the negative community using $n$ tests by testing each community as a whole. In either case, $O(n)$ tests is sufficient.
The above examples show that the lower bound does not necessarily obtained by simple properties of the graph, such as $\mu$ and entropy. Finding a tight lower bound is an interesting future direction.

### v2

@@ 4 Analysis
Let $z=|e^{*}|$. Then we know $\mathbb{E}[{z}]=\mu$ and by Markov inequality, $Pr[z>\mu/\epsilon]<\epsilon$, meaning that with probability $1-\epsilon$, the size of the target edge is at most $\mu/\epsilon$. To exploit this fact, we can modify Algorithm 1 similar to that in the proof of Corollary 4.10. The only difference would be that we would replace $f_{2}(n)$ with $\mu/\epsilon$ in the modification. Thus, $f_{2}(n)$ should be replaced with $\mu/\epsilon$ in the upper bound on the expected number of tests given in the statement of the Corollary.
Replacing $f_{2}(n)$ with $\mu/\epsilon$ it follows from Corollary 4.10 that the expected number of tests is at most $\frac{2}{\log\frac{1}{(1-c)}}H(X)+\frac{2\mu}{\epsilon}$ and the target edge is returned with probability $1-\epsilon.$ Because $c\leq 1/3$, the upper bound on the expected number of tests is minimized when $c=1/3$ and in this case, the expected number of tests is at most $\frac{2}{\log\frac{3}{2}}H(X)+\frac{2\mu}{\epsilon}$. ∎
@@
Markov bound works for general distributions, but it might happen that the mass in $\mathcal{D}$ is concentrated around the mean. In this case, the dependency on $1/\epsilon$ might be slower. For example, if the concentration is exponential, i.e. $Pr(|e^{*}|>2\mu)<2^{-\Omega(n)}$, then with probability $1-2^{-\Omega(n)}$ the algorithm recovers $e^{*}$ and the expected number of tests is at most $\mathbb{E}[L_{1}+L_{2}]\leq\frac{2}{\log\frac{3}{2}}H(X)+2\mu$ (by replacing $f_{2}(n)$ with $2\mu$ in the statement and proof of Corollary 4.10). More precisely, we have the following Corollary.
@@
Suppose $e^{*}$ is sampled from the probability distribution $\mathcal{D}$ and we have $Pr(|e^{*}|>u)=o(1)$. Then, with probability $1-o(1)$, the Algorithm 1 as modified in the proof of Corollary 4.10 (with $f_{2}(n)$ replaced by $u$) returns the correct edge $e^{*}$ using $O(H(X)+u)$ tests with probability $1-o(1)$.
Note that in the above corollary, $u$ can be any stochastic upper bound on the size of the target edge. Thus, the modification of Algorithm 1 does not need to know $\mu$, it just needs to know any stochastic upper bound on the size of the target edge; the guarantee on the expected number of tests however improves as $u$ becomes smaller. Since the smallest value of a stochastic upper bound can be $O(\mu)$, the best guarantee via the above corollary is obtained if $\mu$ is known.
@@ 5 Recovering/Improving Prior Results
In [NSG+21a], as we discussed briefly in Section 2.2, the authors consider $F$ families where each family $j$ has size $M_{j}$, so $n=\sum_{j}M_{j}$. Each family $j$ is infected with probability $q$ and when a family is infected, each of its nodes is infected with probability $p_{j}$ independent of the rest. If a family is not infected, then none of its nodes is infected. They show that to have a constant probability of recovery, one needs $L\geq H(X)=Fh_{2}(q)+\sum_{j=1}^{F}qM_{j}h_{2}\left(p_{j}\right)-w_{j}h_{2}\left(\frac{1-q}{w_{j}}\right)$ where $w_{j}=1-q+q\left(1-p_{j}\right)^{M_{j}}$ and $h_{2}$ is the binary entropy. For the symmetric case where $\forall jp_{j}=p,M_{j}=M$ and hence $w_{j}=w$, the entropy simplifies to $H(X)=Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})$. They provide an algorithm that needs at most $O(Fq(\log(F)+M)+nqp\cdot\log n)$ tests in expectation.
Note that $\mu=nqp$, and by Theorem 4.11, with probability $1-\epsilon$ the target edge can be found using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+\frac{nqp}{\epsilon})$ tests for a constant $c$. If $F\gg 1$ and $M\gg 1$, then the expected number of infections is concentrated around its mean, and with probability $1-o(1)$ the algorithm succeeds using $O(Fh_{2}(q)+qnh_{2}(p)-Fwh_{2}(\frac{1-q}{w})+{nqp})=O(F(h_{2}(q)-wh_{2}(\frac{1-q}{w}))+nq(h_{2}(p)+p))$ tests by Corollary 4.12. This significantly improves the upper bound provided in [NSG+21a] as $Fq(\log(F)+M)\gg F(h_{2}(q)-wh_{2}(\frac{1-q}{w}))$, when fixing $q$ and $F\rightarrow\infty$, and $nqp\cdot\log n\gg nq(h_{2}(p)+p)$ when $n\rightarrow\infty$.
@@
In [ACÖ23], the authors consider a stochastic block infection model (SBIM) where there are $m$ communities and each one has $k$ nodes, so $n=mk$. Each node is a seed with probability $p$ independently. Then every seed node infects nodes in the same community with probability $q_{1}$ and $q_{2}$ for the nodes outside of the community, $q_{1}>q_{2}$. The authors show lower and upper bounds on the number of tests. For the special case where $q_{2}=0$, they further show optimality of the proposed algorithm. The authors show that their algorithm is order-optimal when, $q_{2}=0$, $q=q_{1}$, $kp\ll 1$, $q\ll\frac{1}{\sqrt{k\log(\frac{1}{kp})}}$, $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$.
In Appendix B, we show that with only assuming $kp\ll m^{-\beta}$, the conditions of Corollary 4.12 hold and, the infected set is recovered with $O(H(X)+\mu)$ tests in expectation. In [ACÖ23], the authors prove that when $q_{2}=0$ and $q=q_{1}$, $\mu\leq m\cdot k^{2}\cdot p\cdot(1/k+q)$, and when $kp\ll 1\text{, and }q\ll\frac{1}{\sqrt{k\log\left(\frac{1}{kp}\right)}},$ the entropy is given by $H(X)=m\cdot k\cdot p\cdot\log\left(\frac{1}{kp}\right)+m\cdot k^{2}\cdot p\cdot q\cdot\left(\log k+\log\log\left(\frac{1}{kp}\right)\right)+1$. When $1/k\leq q$, which is a relaxed condition compared to $1\ll kq$, we have $\mu\leq 2m\cdot k^{2}\cdot p\cdot q$. Since $1\ll kq$, it follows that $\mu\leq H(X)$ and hence $O(H(X)+\mu)$ is $O(H(X))$. Now, Corollary 4.12 provides an order-optimal performance guarantee by Theorem 2.2 and the fact that $\epsilon=o(1)$. More generally, [ACÖ23] attains no analytical results, without having the following two conditions: $kp\ll m^{-\beta}\text{ for some fixed $\beta\in(0,1)$, and }1\ll kq$. But, even without these two conditions, we provide the following result due to Theorem 4.11: Algorithm 1 can recover the infected set with probability $1-\epsilon$ using at most $O(H(X)+\mu/\epsilon)$ tests in expectation.
@@ 6 Lower Bound and Optimality of Algorithm 1
In the probabilistic group testing literature, entropy $H(X)$ constitutes an order-wise tight lower bound [LCH+14] (see also Theorem 2.2.) The upper bounds we established in Theorems 4.9-4.11 and Corollary 4.12 involve not only $H(X)$ but also $\mu$, which is the expected number of infections. To shed light on the extent to which our proposed algorithm is order-optimal and our analytical upper bounds are tight, we seek to answer the following questions. (i) Is entropy always a tight lower bound in the generalized group testing problem with general correlation? (ii) Does $\mu$ play a fundamental role (lower-bounding the number of tests) or should one hope to tighten it in our results? (iii) Is the upper bound established in Theorem 4.11 and/or its corresponding algorithm order-wise optimal?
@@ 6.1 When Entropy is a Loose Lower Bound
Graph $G=(V,E)$ has $n$ nodes where each edge of size $n-1$ is included in $E$. To each edge $e$, assign a uniform probability mass $p(e)=\frac{1}{n}$. Figure 2 demonstrates such a graph with $n=4$.
In the above example, the entropy is $\log n$. What makes this example important for us is that any algorithm with a constant probability of success requires at least $\Omega(n)$ tests. This can be proved using a very similar argument as in [GLS22, Theorem 4], and for the sake of completeness, we provide a short proof next. What is significant here is that the gap between the entropy and the required number of tests is $O(n/\log n)$, which is much larger than expected.
Claim 6.2.
Any algorithm with a constant probability of success with the hypergraph in Example 6.1 requires at least $\Omega(n)$ tests.
Proof.
Since one edge is a target edge and each edge consists of $n-1$ nodes, there is only one negative node and the rest of $n-1$ nodes are all positive. Thus, identifying the target edge is equivalent to identifying the negative node. Also, since exactly one node is negative, every test of size two or greater will yield positive results, rendering them redundant. Consequently, the negative node can be identified only by individual testing until the negative node is located. Every node is equally likely to be the negative node. Suppose $k$ tests are performed. Then, the probability that the target negative node is identified is $1-\frac{n-1}{n}\cdot\frac{n-2}{n-1}\cdots\frac{n-k}{n-k+1}=1-\frac{n-k}{n}=k/n$. Thus, $k=\Omega(n)$, for this probability to be constant.
We can generalize the above example as follows:
@@
Graph $G=(V,E)$ has $n$ nodes where at least $r$ fraction of all edges of size $d$ is included in $E$, for a constant $r\in(0,1]$.
The proof of Claim 6.2 can be slightly generalized to prove the following:
Claim 6.4.
There is a distribution over the edges of Example 6.3 such that $H(X)$ is $o(d)$, specifically $O(\log d),$ and any algorithm performing on this distribution, achieving a constant probability of success, needs to perform $\Omega(d)$ tests.
This again shows that the entropy is not a tight lower bound. We will extensively use Example 6.3, Claim 6.4 and its proof presented below to obtain several other lower bounds on the number of tests in this section.
Proof.
Consider a subset $S$ with $d+1$ nodes. It has $d+1$ subsets of nodes of size $d.$ On an average $r$ fraction of these belongs to $E.$ Thus the average number $E(S)$ is $r(d+1).$
Now let the distribution $\mathcal{D}$ be the uniform distribution over the edges in $S$ and 0 elsewhere. Thus, no algorithm needs to test nodes that are not in edges fully in $S.$ Next, for every edge in $S$, there are $d$ nodes in it, and there is exactly one node that is in $S$ but not in the edge; the edge is the target edge if that node is negative and all $d$ nodes in the edge are positive, the $d$ nodes in the edge cover rest of the nodes in $S$. Thus since there is exactly one target edge in $S$, there is exactly one node in $S$ that is negative and the rest of the nodes in $S$ are positive. The number of such potentially negative nodes equals the number of the edges in $S$, which equals $r(d+1)$ in expectation. Determining which edge is the target edge is equivalent to determining which of the candidate nodes is negative. Thus, an algorithm needs to only test the potentially negative nodes. Since only one of the candidate nodes is negative, testing two or more candidate nodes will give a positive outcome. Thus, the candidate nodes must be tested individually to determine which one is negative. Each of the candidate nodes is equally likely to be negative. Thus, arguing as in the proof of Claim 6.2, any algorithm needs to perform $\Omega(m)$ tests for a constant probability of success, where $m$ is the number of potentially negative nodes. Note that $m$ is $r(d+1)$ in expectation. Considering $c$ as a constant, any algorithm needs to perform $\Omega(d)$ tests in expectation for a constant probability of success. Also, $H(X)=O(\log d)$. ∎
Before proceeding further, suppose we run Algorithm 1 on the above examples. It is worth noting that the algorithm bypasses the initial stage and proceeds to test nodes individually in line 1, requiring $n$ tests in Example 6.1 when $0.51/n<c<$ 0.5  [^2] To see this note that in Example 6.1 each node belongs in all but one edge. Thus for each $v$, $w(V\setminus\{v\})$ is the weight of the only edge containing $v$ which is $1/n$. Thus, $w(V\setminus\{v\})<c<1-c$ for each $v.$ And, the algorithm does not satisfy conditions in lines 11, but satisfies the condition in line 1 with $S=V$ for each $v.$ Thus, an empty set is tested in line 1, and testing an empty set is considered negative. Thus, every node is tested in line 1. and $d+2$ tests in Example 6.3 when $0.51/d<c<$ 0.5 and under the distribution mentioned in the the proof of Claim 6.4333Consider the probability distribution in the proof of Claim 6.4) and $0.51/d<c<$ 0.5 . We now argue that Algorithm 1 on example 6.3 bypasses the initial stage and proceeds to test nodes individually in line 1, requiring $d+2$ tests overall. To see this note that each edge in $S$ has $d$ nodes, $S$ has $d+1$ nodes, thus each node in $S$ belongs in all but one edge in $S.$ Also, $q_{v}=0$ for each node outside $S.$ Thus for each $v\in S$, $w(V\setminus\{v\})$ is the weight of the only edge containing $v$ which is $1/d$. Thus, $w(V\setminus\{v\})<c<1-c$ for each $v\in S$. Also, $w(V\setminus\{v\})=1>1-c$ for each $v\in V\setminus S.$ And, the algorithm does not satisfy conditions in line 1 with any $v.$ It removes all nodes in $V\setminus S$ from $V$ in line 1, and $S$ consists of the remaining nodes. Now, the condition in line 1 is satisfied and the nodes outside $S$ are tested together. These nodes test negative. Thus, every node in $S$ is tested in line 1. Thus, there are $d+2$ tests in all.. These examples highlight scenarios where conducting individual tests is optimal, leaving no room for improving the theoretical guarantees in Theorem 4.11 and the following corollaries.
These examples raise the question of finding other lower bounds or identifying classes of graphs where entropy is not a tight lower bound. We generalize Example 6.3 for a larger family of graphs where $H(X)$ is a loose lower bound and at least $\mu$ tests are needed.
@@
Consider $n=|V|$, and let $m$, $d=|e|$ be given parameters. We build a graph with $n$ nodes and $|E|=m+O(d)$ edges where each edge has size $d$. Let the distribution $\mathcal{D}$ be uniform over the edges. Here, $H(X)=\log(m+O(d))$ and $\mu=d$. The graph is constructed in the following manner, where we start with an empty set for the edges and gradually add edges:
@@
The first time that $|E|$ exceeds $m$, $|E|\leq m-1+d+1=m+d.$ We claim that at least $\Omega(d)$ tests are needed to recover the infected set. Suppose $e^{*}$ is the sampled edge, so at some point in line 2 $e^{*}$ was added to $E$. Hence there is a set $S$ of size $d+1$ that contains $e^{*}$ and there are $d$ edges inside of $S$. Now even if the algorithm knows the set $S$, using a similar argument as in [GLS22] and proof of Claim 6.4 for Example 6.3, the algorithm still needs $\Omega(d)$ tests as only one node is negative and it is equally likely to be any of them.
Now if $\mu=d>\log(m+d)\geq H(X)$, then the entropy lower bound is loose and $\mu$ is an order-wise tight lower bound. This happens when $2^{d}>m+d$.
@@ 6.2 When $\mu$ is not a Lower Bound
Consider a graph $G=(V,E)$ comprising $n$ nodes and the edges are of the form $e_{i}=\{v_{1},\ldots,v_{i}\}$ and $p_{e_{i}}=1/n$ for each $1\leq i\leq n$.
@@
Consider a graph with vertex set $V=V_{1}\cup V_{2}\cdots\cup V_{n}$ where $|V_{i}|=n$, so $|V|=n^{2}$ and we call each $V_{i}$ a community. Community $V_{i}$ is composed of nodes $V_{i,j}$ for $j=1,\ldots,n$. The edge set is composed of “small edges” and “large edges”. The set of small edges is defined by $E_{1}=\{e\mid\forall j,k:e=V_{j}\setminus\{V_{j,k}\}\}$. So each small edge is contained in one of the $V_{i}$s and is obtained by removing exactly one node from a $V_{i}$. Thus, each $V_{i}$ has $n$ small edges in it, each of size $n-1$, there are therefore $n^{2}$ small edges. The set of large edges is defined by $E_{2}=\{e\mid\forall i:e=V\setminus V_{i}\}$. Thus, each large edge is obtained by removing all nodes in one $V_{i}$ from $V.$ Thus, each large edge has $n^{2}-n=n(n-1)$ nodes in it, and there are $n$ large edges. Here, $E=E_{1}\cup E_{2}$. The distribution $\mathcal{D}$ puts half of the mass on each of $E_{1},E_{2}$, and the edges in $E_{1}$ (respectively $E_{2}$) have the same probability among themselves. Figure 6 shows an example with $n=4$.
Figure 6: The graph in Example 6.7 with $n=4$, 16 nodes, and 20 edges. Two of the“Small” edges are shown in gray, one contains $\{1,2,3\}$, and the other contains $\{1,3,4\}$. An example of “large” edges is give with red border that contains $\{5,6,\ldots,15,16\}$. There are 16 small edges and 4 large edges in total.
In the above example, $\mu=\Theta(n^{2})$ and $H(X)=O(\log n)$. Following the proof of Claim 6.4 for Example 6.3, it is not hard to see that $\Omega(n)$ tests are needed, because even after knowing whether the sampled edge is in $E_{1}$ or $E_{2}$, the algorithm needs to find the single negative node (in some community $V_{i}$) to recover the sampled edge. Thus, $H(X)$ is a loose lower bound. We next show that $O(n)$ tests are sufficient showing that $\mu$ is not a lower bound at all.
To see that $O(n)$ tests are also sufficient, note that if the infected edge is in $E_{2}$, all nodes except those in one $V_{i}$ will be positive, and no node in that $V_{i}$ is positive. So, one can pick a single node from each $V_{1},V_{2},V_{3},\ldots$ and test the candidates individually. If at least two of them are negative, it means the sampled edge is not in $E_{2}$, otherwise, it is in $E_{2}$. If the sampled edge is not in $E_{2}$, it is in $E_{1}$. This phase consumes at most $n$ tests. If the sampled edge is in $E_{1}$, then all nodes are negative, except for those in one $V_{i}$, and only one node is negative in the exceptional $V_{i}$ and the target edge is in the exceptional $V_{i}$. The exception can be determined by identifying $2$ random nodes in each $V_{i}$, and testing each such pair as a group. Only the pair for the exception will test positive, all others would test negative. Thus, the exception can be identified in additional $n$ tests. Now, the target edge can be identified by testing each node in the exception individually until the negative node is located, and all nodes in the exception except the negative node constitutes the target edge. This requires an additional $n$ tests. Thus, overall, if the target edge is in $E_{1}$, it can be identified in at most $3n$ tests. If the target edge is in $E_{2}$, we can find the negative community using at most $n$ additional tests by individually testing one randomly selected node from each community. Each such test will be positive except for the candidate node from the negative community. Thus, if the target edge is in $E_{2}$, it can be identified in at most $2n$ tests. In either case, $O(n)$ tests is sufficient.
The above examples show that the lower bound can not necessarily be obtained by simple properties of the graph, such as $\mu$ and entropy. Finding a tight lower bound constitutes a future direction.